OpenStudy (anonymous):

choose the equation below that represents the line passing through the point (1,-4) with a slope of 1/2

OpenStudy (harsimran_hs4):

given a point (a, b) and slope m y - b = m (x - a) is the required equation

OpenStudy (anonymous):

ok so can you help me solve it

OpenStudy (harsimran_hs4):

(a, b) = (1, -4) and m = 1/2 so put these values in the above equation and tell what you get?

OpenStudy (anonymous):

so i got to here now im stuck i don't know if its right? y=1/2x + 3?

OpenStudy (harsimran_hs4):

well i think something`s wrong can you show your working ?

OpenStudy (anonymous):

ok y- b =m(x-a) m = 1/2 points (1,-4) y-4= 1/2(x-1) +4 +1 ------------- y = 1/2x + -3?

OpenStudy (harsimran_hs4):

you did it wrong from substituting the values y - (-4) = (1/2)(x-1) y + 4 = (1/2) (x-1) for simplification multiplying both sides by 2 2(y+4) = x-1 2y + 8 = x-1 0 = x-1 - 2y -8 x-2y - 9 = 0 is required equation

OpenStudy (harsimran_hs4):

some times you want it in this form y = 1/2(x-9) (i.e y = mx + c)